# A Low-Rank ADMM Splitting Approach for Semidefinite Programming

### v2

@@ 5.1 Implementation Details
We implement LoRADS [^2] LoRADS is avaliable at https://github.com/COPT-Public/LoRADS in ANSI C with open-source numerical linear algebra libraries, Lapack (Anderson et al., 1999) and Blas (Blackford et al., 2002). To solve the subproblems of LRADMM, we use the warm-started conjugate gradient method to solve the linear systems (22) and (23), where $f(UV^{\top})=\big{\langle}C,UV^{\top}\big{\rangle}$. For the dynamic rank adjustment strategy discussed previously, we initiate the rank at $2\log(m)$ and increment the rank by a factor of 1.5 with each rank update. Specially, given the extensive scale of matrix completion problems from MC_10000 to MC_40000, a modified approach is adopted where the initial rank is set at $\log(m)$. Regarding the timing for variable splitting, we switch to Phase II upon meeting a switching tolerance of 1e-3 in the majority of instances tested. However, for all MaxCut problems (all problems in Section 5.3 and problems identified by starting with ’mcp’ in Section 5.5), we apply a more lenient switching tolerance of 1e-2. Our code is capable of automatically selecting the phase-switching tolerance by analyzing the structural characteristics of the problem. The error used as the indicator for terminating Phase I is primal infeasibility, consistent with the final termination indicator outlined in Section 5.2. Beyond these specific settings, our investigation suggests that the algorithm’s effectiveness relies on applying certain techniques and implementation strategies. In the ensuing subsection, we delve into these aspects in detail.
@@ 5.2 Experiment Setting
  • SDNAL+ terminates when $\max\left\{\frac{\|\mathcal{A}(X)-b\|_{2}}{1+\|b\|_{2}},\frac{\|\mathcal{A}^{*}(y)-C\|_{F}}{1+\|C\|_{F}}\right\}<\epsilon,$ where $\mathcal{A}^{*}(y)=\sum_{i=1}^{m}y_{i}A_{i}.$ Note that Sun et al. (2019) define two additional errors. However, both of them are zeros in our setting.
@@
In the following experiments in this section, we set $\epsilon$ to be 1e-5, and the time limit to be 10,000 seconds [^6] The instances failed to be solved in 10,000 seconds are marked with ”t”. The iteration cap for the LRADMM algorithm is set at 5,000. All the experiments are performed on a Macbook Pro with a 12 Core M3 Pro chip and 18 GB RAM. For clarity, this section only reports the results of solving times for each experiment. A report of properly defined and aligned solving errors for all solvers in all the experiments can be found in Appendix B.
@@ 5.4 The Matrix Completion Problem
The Matrix Completion Problem, which involves estimating the unobserved entries of a partially observed matrix, finds wide application in areas such as collaborative filtering, image inpainting, system identification, among others. SDP is a well-known method for addressing matrix completion problems. However, the use of conventional SDP solvers in matrix completion, particularly in large-scale instances, is hindered by computational inefficiencies (Li et al., 2019). Despite these limitations, recent progress in SDP algorithms, highlighted by Wang and Hu (2023), has demonstrated significant improvements over traditional SDP solvers. Additionally, recent studies by Yalçin et al. (2023) have shown that the Burer-Monteiro factorization method is particularly advantageous in scenarios with limited observations and a small number of measurements. Motivated by these efforts, we test our method on matrix completion problems and demonstrate its effectiveness.
@@
We benchmark our algorithm on a set of randomly generated matrix completion problems, adopting the generating process described by Wang and Hu (2023). The results, as depicted in Table 5, demonstrate that LoRADS consistently outperforms other solvers across a range of instance sizes. For instances larger than MC_10000, further evaluations of the compared solvers were not conducted due to time and computation resource limit. Notably, LoRADS successfully solves an exceptionally large-scale instance with $15,694,167$ constraints and a $40000\times 40000$ matrix variable in less than 6 minutes.
@@ 5.5 General SDP Benchmark Problems
${}^{\star}$ 5991 7230 63.21 2491 943 1208 ${}^{\divideontimes}$ 250 250 0.01 0.01 7.57 0.86
${}^{\star}$ 2167 1743 7.80 124 48.31 35.65 ${}^{\divideontimes}$ 250 250 0.01 0.01 5.42 0.88
${}^{\star}$ 570 10470 513.88 t 728 83.36 ${}^{\divideontimes}$ 250 250 0.01 0.02 5.49 0.79
${}^{\star}$ 2167 1743 6.74 51.46 88.84 19.06 ${}^{\divideontimes}$ 250 250 0.02 0.04 5.25 0.90
${}^{\star}$ 3971 3971 0.95 40.45 6201 20.3 ${}^{\divideontimes}$ 500 500 0.01 0.03 21.36 0.20
${}^{\star}$ 2001 2001 19.89 59.95 4645 12.00 ${}^{\divideontimes}$ 500 500 0.02 0.04 14.59 0.20
${}^{\star}$ 3001 3001 17.21 25.45 3019 17.50 ${}^{\divideontimes}$ 500 500 0.03 0.07 21.87 0.29
${}^{\divideontimes}$ 100 101 0.05 0.14 3.76 0.14 ${}^{\divideontimes}$ 500 500 0.04 0.11 23.22 0.40
${}^{\divideontimes}$ 124 125 0.04 0.26 6.14 0.24 ${}^{\star}$ 2047 1743 24.77 19.03 10.81 18.37
${}^{\divideontimes}$ 124 125 0.07 0.11 4.99 0.23 ${}^{\star}$ 3163 2964 9.57 177 55.38 63.81
${}^{\divideontimes}$ 124 125 0.06 0.10 7.42 0.26 ${}^{\star}$ 9116 9115 4.42 83.95 t 289
${}^{\divideontimes}$ 124 125 0.11 0.32 4.57 0.23 ${}^{\divideontimes}$ 26 136 0.18 0.09 0.48 0.01
${}^{\divideontimes}$ 250 250 0.15 0.71 27.93 1.79 ${}^{\divideontimes}$ 37 229 0.63 1.90 1.69 0.03
${}^{\divideontimes}$ 250 250 0.14 0.49 14.52 1.63 ${}^{\divideontimes}$ 50 358 1.02 2.18 1.65 0.06
${}^{\divideontimes}$ 250 250 0.16 1.59 16.46 2.03 ${}^{\divideontimes}$ 65 529 1.75 4.33 8.89 0.15
${}^{\divideontimes}$ 250 250 0.20 0.61 13.94 1.84 ${}^{\divideontimes}$ 82 748 2.50 4.89 5.34 0.26
${}^{\divideontimes}$ 500 501 0.71 1.84 31.15 1.08 ${}^{\divideontimes}$ 101 1021 2.97 10.19 6.14 0.43
${}^{\divideontimes}$ 500 501 0.72 1.02 106 0.76 ${}^{\star}$ 4029 3540 94.13 73.41 121 12.3
${}^{\divideontimes}$ 500 501 0.62 2.64 35.98 0.67 ${}^{\star}$ 800 4962 53.59 70.77 3068 115
${}^{\divideontimes}$ 500 501 0.65 1.30 36.37 0.76 ${}^{\divideontimes}$ 50 104 0.22 0.32 0.32 0.02
${}^{\star}$ 3163 2964 10.12 199 295 101 ${}^{\divideontimes}$ 100 498 0.57 0.72 1.65 0.10
${}^{\star}$ 1297 1297 116 24.22 628 92.76 ${}^{\divideontimes}$ 150 1106 1.49 2.29 1.57 0.30
${}^{\star}$ 8114 8113 4.91 41.97 t 269 ${}^{\divideontimes}$ 200 1949 3.09 6.01 3.74 0.82
${}^{\divideontimes}$ 100 100 $<$0.005 $<$0.005 1.19 0.07 ${}^{\divideontimes}$ 250 3028 4.67 9.04 3.05 1.89
${}^{\divideontimes}$ 124 124 $<$0.005 0.01 2.05 0.12 ${}^{\divideontimes}$ 300 4375 6.77 19.49 4.22 3.90
${}^{\divideontimes}$ 124 124 $<$0.005 $<$0.005 1.90 0.13 ${}^{\star}$ 601 17979 86.23 127 11.69 9.85
${}^{\divideontimes}$ 124 124 0.01 0.01 1.83 0.12 ${}^{\star}$ 501 37467 71.44 141 6.36 4.57
${}^{\divideontimes}$ 124 124 0.01 0.01 1.56 0.13 ${}^{\star}$ 601 90020 206 369 10.60 8.92
@@
    Problems from SDPLIB are marked with "${}^{\divideontimes}$", and problems from Hans Mittelmann’s SDP benchmarks are marked with "${}^{\star}$"
@@ Acknowledgments
The authors thank Wenzhi Gao from Standford University and Mingyu Lei and Jiayu Zhang from Shanghai Jiao Tong University for their insightful discussions and assistance.

### v3

@@ 5.1 Implementation Details
We implement LoRADS [^2] LoRADS is available at https://github.com/COPT-Public/LoRADS in ANSI C with open-source numerical linear algebra libraries, Lapack (Anderson et al., 1999) and Blas (Blackford et al., 2002). To solve the subproblems of LRADMM, we use the warm-started conjugate gradient method to solve the linear systems (22) and (23), where $f(UV^{\top})=\big{\langle}C,UV^{\top}\big{\rangle}$. For the dynamic rank adjustment strategy discussed previously, we initiate the rank at $2\log(m)$ and increment the rank by a factor of 1.5 with each rank update. Especially, given the extensive scale of matrix completion problems from MC_10000 to MC_40000, a modified approach is adopted where the initial rank is set at $\log(m)$. Regarding the timing for variable splitting, we switch to Phase II upon meeting a switching tolerance of 1e-3 in the majority of instances tested. However, for all MaxCut problems (all problems in Section 5.3 and problems identified by starting with ’mcp’ in Section 5.5), we apply a more lenient switching tolerance of 1e-2. Our code is capable of automatically selecting the phase-switching tolerance by analyzing the structural characteristics of the problem. The error used as the indicator for terminating Phase I is primal infeasibility, consistent with the final termination indicator outlined in Section 5.2. Beyond these specific settings, our investigation suggests that the algorithm’s effectiveness relies on applying certain techniques and implementation strategies. In the ensuing subsection, we delve into these aspects in detail.
@@ 5.2 Experiment Setting
  • Denote the slack variable of the dual of (1) by $S$ and the metric projection of $(X-S)$ onto $S_{+}^{n}$ by $\Pi_{S_{+}^{n}}(X-S)$, SDPNAL+ terminates when $\max\left\{\frac{\|\mathcal{A}(X)-b\|}{1+\|b\|},\frac{\left\|\mathcal{A}^{*}(y)+S-C\right\|}{1+\|C\|},\frac{1}{5}\frac{\left\|X-\Pi_{S_{+}^{n}}(X-S)\right\|}{1+\|X\|+\|S\|}\right\}\leq\epsilon$.
@@
In the following experiments in this section, we set $\epsilon$ to be 1e-5, and the time limit to be 10,000 seconds [^6] The instances that failed to be solved in 10,000 seconds are marked with ”t”. The iteration cap for the LRADMM algorithm is set at 5,000. All the experiments are performed on a Macbook Pro with a 12 Core M3 Pro chip and 18 GB RAM. For clarity, this section only reports the results of solving times for each experiment. A report of properly defined and aligned solving errors for all solvers in all the experiments can be found in Appendix B.
@@ 5.4 The Matrix Completion Problem
The Matrix Completion Problem, which involves estimating the unobserved entries of a partially observed matrix, finds wide application in areas such as collaborative filtering, image inpainting, and system identification, among others. SDP is a well-known method for addressing matrix completion problems. However, the use of conventional SDP solvers in matrix completion, particularly in large-scale instances, is hindered by computational inefficiencies (Li et al., 2019). Despite these limitations, recent progress in SDP algorithms highlighted by Wang and Hu (2023), has demonstrated significant improvements over traditional SDP solvers. Additionally, recent studies by Yalçin et al. (2023) have shown that the Burer-Monteiro factorization method is particularly advantageous in scenarios with limited observations and a small number of measurements. Motivated by these efforts, we test our method on matrix completion problems and demonstrate its effectiveness.
@@
We benchmark our algorithm on a set of randomly generated matrix completion problems, adopting the generating process described by Wang and Hu (2023). The results, as depicted in Table 5, demonstrate that LoRADS consistently outperforms other solvers across a range of instance sizes. For instances larger than MC_10000, further evaluations of the compared solvers were not conducted due to time and computation resource limits. Notably, LoRADS successfully solves an exceptionally large-scale instance with $15,694,167$ constraints and a $40000\times 40000$ matrix variable in less than 6 minutes.
@@ 5.5 General SDP Benchmark Problems
AlH 5991 7230 63.21 2491 943 1208 mcp250-1 250 250 0.01 0.01 7.57 0.86
BH2 2167 1743 7.80 124 48.31 35.65 mcp250-2 250 250 0.01 0.01 5.42 0.88
cancer_100 570 10470 513.88 t 728 83.36 mcp250-3 250 250 0.01 0.02 5.49 0.79
CH2 2167 1743 6.74 51.46 88.84 19.06 mcp250-4 250 250 0.02 0.04 5.25 0.90
checker_1.5 3971 3971 0.95 40.45 6201 20.3 mcp500-1 500 500 0.01 0.03 21.36 0.20
G40_mb 2001 2001 19.89 59.95 4645 12.00 mcp500-2 500 500 0.02 0.04 14.59 0.20
G48_mb 3001 3001 17.21 25.45 3019 17.50 mcp500-3 500 500 0.03 0.07 21.87 0.29
gpp100 100 101 0.05 0.14 3.76 0.14 mcp500-4 500 500 0.04 0.11 23.22 0.40
gpp124-1 124 125 0.04 0.26 6.14 0.24 NH2 2047 1743 24.77 19.03 10.81 18.37
gpp124-2 124 125 0.07 0.11 4.99 0.23 NH3 3163 2964 9.57 177 55.38 63.81
gpp124-3 124 125 0.06 0.10 7.42 0.26 p_auss2_3.0 9116 9115 4.42 83.95 t 289
gpp124-4 124 125 0.11 0.32 4.57 0.23 qap5 26 136 0.18 0.09 0.48 0.01
gpp250-1 250 250 0.15 0.71 27.93 1.79 qap6 37 229 0.63 1.90 1.69 0.03
gpp250-2 250 250 0.14 0.49 14.52 1.63 qap7 50 358 1.02 2.18 1.65 0.06
gpp250-3 250 250 0.16 1.59 16.46 2.03 qap8 65 529 1.75 4.33 8.89 0.15
gpp250-4 250 250 0.20 0.61 13.94 1.84 qap9 82 748 2.50 4.89 5.34 0.26
gpp500-1 500 501 0.71 1.84 31.15 1.08 qap10 101 1021 2.97 10.19 6.14 0.43
gpp500-2 500 501 0.72 1.02 106 0.76 sensor_500 4029 3540 94.13 73.41 121 12.3
gpp500-3 500 501 0.62 2.64 35.98 0.67 shmup4 800 4962 53.59 70.77 3068 115
gpp500-4 500 501 0.65 1.30 36.37 0.76 theta1 50 104 0.22 0.32 0.32 0.02
H3O 3163 2964 10.12 199 295 101 theta2 100 498 0.57 0.72 1.65 0.10
hand 1297 1297 116 24.22 628 92.76 theta3 150 1106 1.49 2.29 1.57 0.30
ice_2.0 8114 8113 4.91 41.97 t 269 theta4 200 1949 3.09 6.01 3.74 0.82
mcp100 100 100 $<$0.005 $<$0.005 1.19 0.07 theta5 250 3028 4.67 9.04 3.05 1.89
mcp124-1 124 124 $<$0.005 0.01 2.05 0.12 theta6 300 4375 6.77 19.49 4.22 3.90
mcp124-2 124 124 $<$0.005 $<$0.005 1.90 0.13 theta12 601 17979 86.23 127 11.69 9.85
mcp124-3 124 124 0.01 0.01 1.83 0.12 theta102 501 37467 71.44 141 6.36 4.57
mcp124-4 124 124 0.01 0.01 1.56 0.13 theta123 601 90020 206 369 10.60 8.92
@@
    Problems from SDPLIB are marked with "", and problems from Hans Mittelmann’s SDP benchmarks are marked with ""
@@ Acknowledgments
The authors thank Wenzhi Gao from Stanford University and Mingyu Lei and Jiayu Zhang from Shanghai Jiao Tong University for their insightful discussions and assistance.